OpenStudy (anonymous):

What is the magnitude of the electric field at a point midway between a -5.8μC and a +6.0μC charge 8.4cm apart? Assume no other charges are nearby.

OpenStudy (aaronq):

Force: F= \(\dfrac {q_1* q_2}{4\pi\epsilon_0r^2}\) Electric Field: |E|= \(\dfrac{F}{q}\)

OpenStudy (jt950):

E=1.4x10^-4 N/C?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."